A sample of aspirin (acetyl salicyclic acid) of molecular mass 180, on analysis was found to contain 60% carbon, 4.4% hydrogen and 35.6% oxygen

i.Determine the empirical formula of aspirin

ii. What is the molecular formula of aspirin?

iii. State the technique that can be used to determine the molecular weight of aspirin


1
Expert's answer
2021-02-12T04:55:03-0500

% of Carbon = 60 % .

% of hydrogen = 4.4 % .

% of Oxygen = 35.6 % .


Moles of carbon = 60 / 12 = 5 mol .

Moles of hydrogen = 4.4 / 1 = 4.4 mol.

Moles of oxygen = 35.6 / 16 = 2.225 mol.


When we devide each by 2.2 and multiply by 4

Moles of C = 5 / 2.2 = 2.25 × 4 = 9

Moles of hydrogen = 4.4 / 2.2 = 2 ×4 = 8

Moles 0f Oxygen = 2.2 / 2.2 =1 × 4 = 4


Emperical Formula = C9H8O4.



Emperical mass = 9 ×12 + 1×8 + 4×16 = 180 .


n = Molecular mass / Emperical mass .


n = 180 / 180 = 1 .


Molecular Formula = C9H8O4.